OpenStudy (anonymous):

The perimeter of a rectangle is 20ft , and the area of the rectangle is 52ft^2 . I need to find the dimensions of the rectangle. Is anyone good at word problems?

OpenStudy (anonymous):

Solve:\[2 \times (a+b)=20\] and \[a \times b=52\] Find a and be, being a and b the dimensions

OpenStudy (anonymous):

ok Im still lost. :(

OpenStudy (anonymous):

what is the difference between the perimeter and the area?

OpenStudy (anonymous):

perimeter is sum of four dimensions, area is product of two different dimensions 2*(a+b) = 20 so a+b = 10 then b = 10 - a a*b = 20 so a(10 - a) = 20 then a^2 - 10a + 20 = 0

OpenStudy (anonymous):

Imagine you have a ranch. And that it is surrounded by a fence. The total length of the fence is the perimeter. The surface occupied by the ranch is the area

OpenStudy (anonymous):

so for the perimeter it would be 6/4?

OpenStudy (anonymous):

2(6) ad 2(4)

OpenStudy (anonymous):

and*